Consider a monopolist whose total cost function is TC = 40 + 4Q + Q2 and whose marginal cost function is MC = 4 + 2Q. The demand function for the firm’s good is P = 160 - 0.5Q. The firm optimizes by producing the level of output that maximizes profit or minimizes loss. If the firm uses a uniform pricing strategy, then rounded to the nearest dollar the deadweight loss that results is:
$27
$135
$284
$412
1
Expert's answer
2014-11-27T12:58:36-0500
TC = 40 + 4Q + Q2, MC = 4 + 2Q, P = 160 - 0.5Q
If the firm optimizes by producing the level of output that maximizes profit or minimizes loss, then output is produced at the point, where MR = MC:
MR = TR' = 160 - Q
160 - Q = 4 + 2Q
Q = 52 units
Level of P at MR = MC is P = 4 + 2*52 = $108
Price of the monopolist is Pm = 160 - 0.5*52 = $134
The price discremination output is at MC = D:
4 + 2Q = 160 - 0.5Q
Q = 62.4
If the firm uses a uniform pricing strategy, then rounded to the nearest dollar the deadweight loss that results is:
DWL = 0.5*(134 - 108)*(62.4 - 52) = 0.5*26*10.4 = $135.2
So, the right answer is b) $135.